OpenStudy (anonymous):

please help. Studying for my exam, thank you. log b 5 = 1.1610 log b 2 = .5000 Find each of the following. (a)log(base b)125 (b) log(base b)3√25

OpenStudy (anonymous):

for a i got 3log (baseb) 5

OpenStudy (campbell_st):

1st one \[\log _{b} 5^3 = 3\log _{b} 5\] the other appears to be \[\log _{b} \sqrt[3]{5} = \log _{b} 5^{\frac{1}{3}} = \frac{1}{3} \log _{b} 5\]

OpenStudy (anonymous):

thanks for your help!

OpenStudy (campbell_st):

hope it helps